Jim can fill a pool carrying buckets of water in 30 minutes. Sue can do the same job in 45 minutes. Tony can do the same job in 1 ½ hours. How quickly can all three fill the pool together
Answer by AnlytcPhil(1116)

Jim can fill a pool carrying buckets of water in 30 minutes. Sue can do the
same job in 45 minutes. Tony can do the same job in 1 ½ hours. How quickly can
all three fill the pool together.


What we need is this equation:
Jim's rate + Sue's rate + Tony's rate = their combined rate

>>...How quickly can all three fill the pool together...<<

Let the time be x minutes.
THEREFORE
their combined rate = 1 pool per x minutes or
their combined rate = 1/x pool per minute

>>...Jim can fill a pool...in 30 minutes...<<
TRANSLATION:
Jim's rate is 1 pool per 30 minutes or
Jim's rate = 1/30 pool per minute

>>...Sue can do the same job in 45 minutes...<<
TRANSLATION:
Sue's rate is 1 pool per 45 minutes or
Sue's rate = 1/45 pool per minute

>>...Tony can do the same job in 1 ½ hours.,,<<
TRANSLATION:
Tony's rate is 1 pool per 1½ hour
Tony's rate is 1 pool per 90 minutes
Tony's rate = 1/90 pool per minute
Now
Jim's rate + Sue's rate + Tony's rate = their combined rate
becomes
1/30 + 1/45 + 1/90 = 1/x
Get the LCD of all the denominators 30, 45, 90, and x, which is 90x.
Multiply through by 90x
(90x)(1/30) + (90x)(1/45) + (90x)(1/90) = (90x)(1/x)
3x + 2x + x = 90
6x = 90
x = 90/6
x = 15 minutes
Edwin